OpenStudy (anonymous):

solve for x: 6x - 4y = 3 please hurry

OpenStudy (alexandervonhumboldt2):

6x=3+4y x=1/2+2/3y

OpenStudy (alexandervonhumboldt2):

plz medal and fan if i'm correct